OpenStudy (anonymous):

find all the solutions of the equation in the interval [0 2pi] 2 sin^2 x= 2 + cos x

OpenStudy (mertsj):

Let's replace sin^2x with 1-cos^2x

OpenStudy (mertsj):

\[2(1-\cos ^2x)=2+\cos x\]

OpenStudy (mertsj):

\[2-2\cos ^2x=2+\cos x\]

OpenStudy (mertsj):

\[-2\cos ^2x-\cos x=0\]

OpenStudy (mertsj):

\[-\cos x(2\cos x+1)=0\]

OpenStudy (mertsj):

Perhaps you can finish it now.

OpenStudy (anonymous):

yes! thanks!!

OpenStudy (mertsj):

yw